MICROWAVE SPECTRAL STUDIES OF BENZYL ALCOHOL AND SEC-PHENETHANOL

\underline{KARISSA UTZAT}, ROBERT K. BOHN, AND EMILY CHU, Department of Chemistry, University of Connecticut, Storrs, CT 06269-3060.

Benzyl compounds (C_{6}H_{5}C_{α}H_{2}X), generally fall into two conformational classes. With X=CH_{3}, C_{2}H_{5}, F, or Cl, the C-C-C_{α}-X dihedral angle is 90^{∘}; with X=CN and CCH, the dihedral angle is 0^{∘}; however, benzyl alcohol (C_{6}H_{5}CH_{2}OH) and benzyl amine (C_{6}H_{5}CH_{2}NH_{2}) have asymmetric structures. Two conformers of benzyl amine have been observed, a one with a dihedral angle approximately 50^{∘} and a higher energy form with dihedral angle 90^{∘}. Benzyl alcohol has a single asymmetric conformation with dihedral, C-C-C_{α}-O about 60^{∘} ^{b}. The microwave rotational spectrum displays tunneling doublets for all transitions indicating a low barrier path between two of the four equivalent energy minima. These splittings are significantly smaller in the C_{6}H_{5}CH_{2}OD spectrum. The rotational spectrum of sec-phenethanol, (C_{6}H_{5}C_{α}H(C_{β}H_{3})OH), displays a single conformation with C-C-C_{α}-C_{β} dihedral angle approximately 90^{∘} and the C-C-C_{α}-O dihedral angle approximately 30^{∘}, and no tunneling split transitions. ab initio calculations on sec-phenethanol and benzyl alcohol will also be discussed.

aS. Melandri, A. Maris, P.G. Favero, W. Caminati, CHEMPHYSCHEM 2(3), 172-177 (2001).

bM. Trætteberg, H. Østensen, R. Seip, Acta Chemica Scandinavica A 34, 449-454 (1980).